OpenStudy (anonymous):

second one:Calculus: Integrals, antiderivatives, and Solving Differential Equations Symbolically:Separation of variables y'=cos^(2)(y)/sin^(2)(x)

OpenStudy (anonymous):

\[\int \frac{dy}{cos^2{y}}=\int \frac{dx}{sin^2{x}}\] Now that it's separated you can integrate.

OpenStudy (mathmate):

dy/dx =cos^2(y)/sin^2(x) dy/cos^2(y) = dx/sin^2(x) sec^2(y)dy = csc^2(x) dx Each side is a standard integral, integrate each side to get the answer. Do not forget the integration constant.

OpenStudy (anonymous):

thank you! this is really helpful :) okay i think i can do the rest of my homework. thanks to you all again!

OpenStudy (earthcitizen):

no p! yw

OpenStudy (anonymous):

no worries, good luck :)